OpenStudy (anonymous):

For -1 < x < 1, by sketching a triangle, find cos^{-1} x + sin^{-1}x = _____ degrees.

OpenStudy (pawanyadav):

Let's check for x=1/2 Cos^{-1} 1/2 + sin ^{-1} 1/2 =π/3+π/6=π/2 And this is true for every value of x from -1 to 1.. So answer is π/2 or 90°

OpenStudy (pawanyadav):

@Kimes